# A two-phase model of galaxy formation: III. The formation of globular clusters

### v1

@@ Abstract
We develop a model of globular cluster (GC) formation within the cosmological hierarchy of structure formation. The model is rooted in the ‘two-phase’ scenario of galaxy formation developed in Paper-I, where the fast accretion of dark matter halos at high redshift leads to the formation of self-gravitating, turbulent gas clouds that subsequently fragment into dynamically hot systems of dense sub-clouds with masses $\sim 10^{6}$$10^{7}M_{\odot}$. Here we elaborate on the formation, evolution, and fate of these sub-clouds, and show that some of the sub-clouds can be compactified via two distinctive channels into a ‘supernova-free’ regime to form two distinct populations of GCs. The model is simple, characterized by a small number of free parameters underpinned by physical considerations, and can be efficiently implemented into cosmological N-body simulations to generate a coherent sample of halos, galaxies, and GCs. Our model can reproduce a range of observations on GCs, including the mass function, the size-mass relation, the frequency per unit host galaxy/halo mass, the bimodal metallicity distribution, and the spatial profile. Predictions for GCs are made for both the local Universe and for redshift up to $z\approx 10$, and can be tested by upcoming observations.
@@ 1 Introduction
Globular clusters (hereafter GCs) are massive and dense stellar systems observed in nearly every galaxy (Harris & Racine, 1979; Krumholz et al., 2019). Their old age, low metallicity and high density, compared to their host galaxies, imply that they are remnants of star formation in the early Universe (Peebles & Dicke, 1968; Fall & Rees, 1985) and thus provide fossils for cosmological archaeology. At a distinct and important rank in the cosmic structure hierarchy, GCs are crucial to understanding galaxy formation, extending its boundaries to extreme conditions on small scales.
@@
Current observations have already provided critical information about the formation of GCs. They are observed in galaxies of all masses (Blakeslee, 1997, 1999; Harris & Racine, 1979; Harris, 2023) and morphology (Harris et al., 2015) in diverse environments (Forbes et al., 2020; Janssens et al., 2022; Saifollahi et al., 2022; Jones et al., 2023; Forbes & Gannon, 2024) in the local Universe. The discovery of young massive star clusters (YMSCs; Whitmore et al., 1999; Figer et al., 1999; Schweizer, 1999; Figer, 2004; Bastian et al., 2006) suggests that the conditions for GC formation are still present in the local Universe, particularly in starburst and interacting/merging galaxies (see Krumholz et al., 2019, for a review).
@@
The ubiquitous presence of GCs implies that their formation may be driven by some general formation mechanisms rather than rare conditions associated with special galaxies or with rare stochastic processes. On the other hand, observations also show that the population of GCs within a galaxy, commonly referred to as the globular cluster system (GCS) of the galaxy, shows strong correlations with the properties of the host galaxy and host halo (Harris & Racine, 1979), suggesting that the diversity and complexity of the GC population may be related to those of the galaxy population. The most intriguing correlation is the linear relationship between the total mass of GCS and the host halo mass observed in the local Universe (e.g. Blakeslee, 1997, 1999; Spitler & Forbes, 2009; Harris et al., 2013; Hudson et al., 2014; Harris et al., 2015, 2017; Dornan & Harris, 2023; Harris, 2023). Exceptions seem to occur in dwarf galaxies with halo masses $\sim 10^{10}{\rm M_{\odot}}$, where the number of GCs is very small and has large scatter (Saifollahi et al., 2022; Jones et al., 2023), and in the brightest cluster galaxies (BCGs), where contributions from satellite galaxies and the intra-cluster population can introduce significant ambiguity (Spitler & Forbes, 2009; Hudson et al., 2014; Dornan & Harris, 2023).
@@
This linear relation was inferred to be a result of the central limit theorem applied to massive galaxies, where the total number of GCs brought in by merger events naturally normalize the GCS mass with the halo mass (Boylan-Kolchin, 2017). However, the establishment of such a relation in dwarf galaxies and whether or not this relation also holds at high redshift (hereafter high-$z$ ) remain poorly understood (e.g. Bastian et al., 2020). Within individual galaxies, GCs are found to have a bimodal distribution in color (Fahrion et al., 2020b; Harris, 2023; Hartman et al., 2023). This bimodality is interpreted as a result of the bimodal distribution of metallicity, as metal-poor and metal-rich stars follow different evolutionary paths in the color-magnitude diagram. The observed bimodal distribution signifies a bimodal formation mechanism. There is thus a strong motivation to investigate how the formation of GCs is regulated by the metallicity of the interstellar medium (ISM).
@@
The main challenge in modeling GC formation in a cosmological context is the large gap between the properties of GCs and their host galaxies. GCs are relatively small, with a typical size of $\sim 10\,{\rm pc}$ (e.g. Krumholz et al., 2019) that is about two orders of magnitude smaller than a typical dwarf galaxy (e.g. Wang et al., 2023). Consequently, certain approximations are necessary to simulate GC formation as a part of galaxy formation in the cosmic density field. Along this line, two semi-analytic approaches have been developed, each incorporating different physical inputs derived from numerical simulations. The first approach involves post-processing cosmological hydrodynamic simulations, or their zoom-in runs, with subgrid physics specific to GC formation. Depending on the resolution of the simulation used, the input conditions for GC formation can range from global properties of individual GMCs (e.g. Grudić et al., 2023) to the smoothed field of the gas properties (e.g. Kruijssen et al., 2011; Pfeffer et al., 2018; Kruijssen et al., 2019a). The second approach more closely resembles traditional semi-analytic models of galaxies based on dark matter halos and their merger trees. The global properties of individual galaxies are obtained from a semi-analytic model (El-Badry et al., 2019; De Lucia et al., 2023) or an empirical model of galaxy formation (Li & Gnedin, 2014; Choksi et al., 2018), while the criteria for GC formation are based on some properties of the assembly histories of galaxies/halos, such as mergers of galaxies (El-Badry et al., 2019) and rates of halo mass assembly (Li & Gnedin, 2014; Choksi et al., 2018; Chen & Gnedin, 2022). Both approaches have been successful in reproducing some observed properties of GCs. Despite the progress, significant knowledge gaps still exist in our understanding of the GC population. Indeed, our understanding of galaxy formation, in particular of star formation and feedback processes, is still incomplete, and so the large-scale conditions for the formation of GCs are still uncertain. In addition, small-scale processes driving the formation and disruption of GCs are also not well understood.
@@
Recent observations by the James Webb Space Telescope (JWST) have opened a new avenue to study galaxy formation in the early Universe. Recent observations indicated that early galaxies may differ significantly from their counterparts in the local Universe, suggesting that the processes driving galaxy formation at high $z$ may be different from those in the low-$z$ universe. In particular, the efficiency of gas assumption by star formation seems to be enhanced significantly at high $z$ (Bouwens et al., 2023; Finkelstein et al., 2023; Xiao et al., 2023; Weibel et al., 2024; Wang et al., 2024), supermassive black holes (SMBHs) appear to be over-massive for their host galaxy mass (Maiolino et al., 2023), there seems to be an inversion of metallicity gradients in both the ISM and circumgalactic medium (CGM; Wang et al., 2022; Lin et al., 2023), and the total mass of bound star clusters in galaxies seems to be excessively large (Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024). Given that many GCs are old, these discoveries at high $z$ have important implications for understanding GC formation.
@@
Motivated partly by these discoveries, Mo et al. (2023, hereafter Paper-I) recently proposed a ‘two-phase’ model of galaxy formation that highlights the transition from an early phase of fast halo accretion to a later phase of slow accretion. A critical part of this model is the formation of self-gravitating and turbulent gas clouds (SGCs) during the fast phase. Such an SGC is expected to fragment subsequently and form a dynamically hot system of dense sub-clouds (SCs), each comparable to a GC in terms of both size and mass. This model is capable of reproducing numerous properties of galaxies observed over a large range of redshift. In particular, its predictions for the sub-cloud formation are directly related to the formation of GCs. As we will show in this paper, the model naturally provides conditions for two channels for the formation of GCs, and the GCs produced in the two channels have distinctive properties. Our model thus provides a unified cosmological framework to understand the formation of both galaxies and their GCs, allowing the model to be tested on spatial scales ranging from star clusters to large-scale structures, and on time scales from $z\approx 0$ to $z\approx 10$.
@@
This paper is organized as follows. In §2, we describe the context of galaxy formation within which GCs emerge. In §3, we elaborate on our model of GC formation. In §4, we describe numerical implementations of our model, present its predictions and compare them with observations and with other models. In §5, we summarize our results and discuss their implications. Throughout this paper, we adopt a flat $\Lambda$CDM cosmology with parameters derived from the Planck2015 results (Ade et al., 2016): the density parameters $\Omega_{\rm M,0}=0.3089$, $\Omega_{\rm B,0}=0.0486$, and $\Omega_{\rm\Lambda,0}=0.6911$; the Hubble constant $H_{0}=100\,h\,{\rm km\,s^{-1}\,Mpc^{-1}}$, with $h=0.6774$; the Gaussian initial density field with a power spectrum $P(k)\propto k^{n}$ with $n=0.9667$; and an amplitude specified by $\sigma_{8}=0.8159$. To avoid confusion, we use ‘log’ to denote the base-10 logarithm, and ‘ln’ to denote the natural logarithm. We use the terms 1, 2, and 3-$\sigma$ ranges to denote those centralized at the median value and covering $68\%$, $95\%$, and $99.7\%$, respectively, of the probability mass for any probability distribution.
@@ 2.1 The assembly of dark matter halos
Our modeling of the halo population is based on a sample of halos and merger trees constructed from numerical simulations of the current $\Lambda$CDM cosmology, as described in §4. For each subhalo merger tree rooted in a ‘root subhalo’ (i.e. the subhalo without a descendant), we decompose it into a set of disjoint branches using the method described in Chen et al. (2023b, see their §3.2.1). This is equivalent to traversing over the tree in a depth-first order starting from the root subhalo, and extracting each subsequent segment of the tree bracketed by a subhalo and its main leaf. For each branch, we find the last snapshot when it is a central subhalo, and define this snapshot as the ‘infall snapshot’. The assembly history at and before this snapshot is defined as the ‘central stage’ of the branch, while that after it is defined as the ‘satellite stage’.
@@
For each branch in the merger tree, we smooth its virial mass history, $M_{\rm v}(z)$, by a running Gaussian kernel with a width equal to the dynamical time scale, $t_{\rm dyn}=R_{\rm v}(z)/V_{\rm v}(z)$, where $R_{\rm v}(z)$ and $V_{\rm v}(z)$ are, respectively, the virial radius and virial velocity of the halo at the redshift $z$ in question. Smoothed versions of other virial quantities are defined using the smoothed $M_{\rm v}(z)$, while histories of other relevant halo quantities are smoothed using the same window as that used for $M_{\rm v}(z)$. The smoothing is required to ensure the numerical stability of our model implementation.
@@ 2.2 The collapse of the gas component
where the cosmic baryon number density $n_{\rm B,0}=3H_{0}^{2}f_{\rm B}\Omega_{\rm m}/(8\pi Gm_{\rm p})\approx 2.48\times 10^{-7}\,{\rm cm}^{-3}$, $\Delta_{\rm v}=200$, $f_{\rm gas}=0.04f_{\rm gas,0.04}$, $\mu=1.2$. The factor $f_{\rm str}$ characterizes the ‘streamness’ of the gas inflow. For spherically symmetric contraction, $f_{\rm str}=1$, and the typical value of $f_{\rm gas}$, $0.04$, is regulated by feedback (see Chen et al., 2023a, hereafter Paper-II). For cold-mode streams, $1<f_{\rm str}\leq f_{\rm B}/f_{\rm gas}\approx 4$, as gas can penetrate into a galaxy along filaments and a larger fraction of gas can pile up around the halo center. Mandelker et al. (2018) and Dekel et al. (2023) estimated that the effective radius of streams is about $0.043R_{\rm v}$, similar to the contraction factor based on $f_{\rm gas}$. The SGC density given by Eq. (2.2) is much lower than the supernova-free threshold $n_{\rm snf}\approx 10^{3.5}\,{\rm cm}^{-3}$ required to form GCs (see §3.5) at $z=2$, and marginally touches the threshold at $z=9$. Thus, additional processes within SGCs are required for the formation of GCs in galaxies.

### v2

@@ Abstract
We develop a model of globular cluster (GC) formation within the cosmological hierarchy of structure formation. The model is rooted in the ‘two-phase’ scenario of galaxy formation developed in Paper-I, where the fast accretion of dark matter halos at high redshift leads to the formation of self-gravitating, turbulent gas clouds that subsequently fragment into dynamically hot systems of dense sub-clouds with masses $\sim 10^{6}$$10^{7}{\rm M_{\odot}}$. Here we elaborate on the formation, evolution, and fate of these sub-clouds, and show that some of the sub-clouds can be compactified via two distinctive channels into a ‘supernova-free’ regime to form two distinct populations of GCs. The model is simple, characterized by a small number of free parameters underpinned by physical considerations, and can be efficiently implemented into cosmological N-body simulations to generate a coherent sample of halos, galaxies, and GCs. Our model can reproduce a range of observations on GCs, including the mass function, the size-mass relation, the frequency per unit host galaxy/halo mass, the bimodal metallicity distribution, and the spatial profile. Predictions for GCs are made for both the local Universe and for redshift up to $z\approx 10$, and can be tested by upcoming observations.
@@ 1 Introduction
Globular clusters (hereafter GCs) are massive and dense stellar systems observed in nearly every galaxy (Harris & Racine, 1979; Krumholz et al., 2019). Many GCs have old age, low metallicity and high density compared to their host galaxies, implying that they are remnants of star formation in the early Universe (Peebles & Dicke, 1968; Fall & Rees, 1985) and thus provide fossils for cosmological archaeology. At a distinct and important rank in the cosmic structure hierarchy, GCs are crucial to understanding galaxy formation, extending its boundaries to extreme conditions on small scales.
@@
Current observations have already provided critical information about the formation of GCs. They are observed in galaxies of all masses (Blakeslee, 1997, 1999; Harris & Racine, 1979; Harris, 2023) and morphology (Harris et al., 2015) in diverse environments (Forbes et al., 2020; Janssens et al., 2022; Saifollahi et al., 2022; Jones et al., 2023; Forbes & Gannon, 2024) in the local Universe. The discovery of young massive star clusters (YMSCs; Whitmore et al., 1999; Figer et al., 1999; Schweizer, 1999; Zhang & Fall, 1999; Figer, 2004; Bastian et al., 2006) suggests that the conditions for GC formation are still present in the local Universe, particularly in starburst and interacting/merging galaxies (see Krumholz et al., 2019, for a review).
@@
The ubiquitous presence of GCs implies that their formation may be driven by some general formation mechanisms rather than rare conditions associated with special galaxies or with rare stochastic processes. On the other hand, observations also show that the population of GCs within a galaxy, commonly referred to as the globular cluster system (GCS) of the galaxy, shows strong correlations with the properties of the host galaxy and host halo (Harris & Racine, 1979), suggesting that the diversity and complexity of the GC population may be related to those of the galaxy population. The most intriguing correlation is the linear relationship between the total mass of GCS and the host halo mass observed in the local Universe (e.g. Blakeslee, 1997, 1999; Spitler & Forbes, 2009; Harris et al., 2013; Hudson et al., 2014; Harris et al., 2015, 2017; Dornan & Harris, 2023; Harris, 2023). Exceptions seem to occur in dwarf galaxies with halo masses $\sim 10^{10}{\rm M_{\odot}}$, where the number of GCs is very small and has large scatter (Saifollahi et al., 2022; Jones et al., 2023), and in the brightest cluster galaxies (BCGs), where contributions from satellite galaxies and the intracluster population can introduce significant ambiguity (Spitler & Forbes, 2009; Hudson et al., 2014; Dornan & Harris, 2023).
@@
This linear relation was inferred to be a result of the central limit theorem applied to massive galaxies, where the total number of GCs brought in by merger events naturally normalize the GCS mass with the halo mass (Boylan-Kolchin, 2017; Valenzuela et al., 2021). However, the establishment of such a relation in dwarf galaxies and whether or not this relation also holds at high redshift (hereafter high $z$) remain poorly understood (e.g. Bastian et al., 2020). Alternative explanations have also been proposed, such as higher gas fractions and enhanced GC formation rates in halos of lower masses (e.g. Choksi et al., 2018; Choksi & Gnedin, 2019; El-Badry et al., 2019; Chen & Gnedin, 2022, 2023, 2024). Within individual galaxies, GCs are found to have a bimodal distribution in color (Fahrion et al., 2020a; Harris, 2023; Hartman et al., 2023). This bimodality is interpreted as a result of the bimodal distribution of metallicity, as metal-poor and metal-rich stars follow different evolutionary paths in the color-magnitude diagram, which is also supported by observations with spectroscopically calibrated color-metallicity conversions (Fahrion et al., 2020b; Hartman & Harris, 2024). The observed bimodal distribution signifies mechanisms that correlate non-linearly with metallicity, such as galaxy-galaxy merger (Li & Gnedin, 2014; Choksi et al., 2018; Valenzuela et al., 2021) and high-density/pressure condition (Kruijssen, 2012; Ma et al., 2020), or a bimodal pair of mechanisms that drives the formation and evolution of GCs. There is thus a strong motivation to investigate how the formation and evolution of GCs are correlated with the metallicity of the interstellar medium (ISM).
@@
The main challenge in modeling GC formation in a cosmological context is the large gap between the properties of GCs and their host galaxies. GCs are relatively small, with a typical size of $\sim 10\,{\rm pc}$ (e.g. Krumholz et al., 2019) that is about two orders of magnitude smaller than a typical dwarf galaxy (e.g. Wang et al., 2023). Consequently, certain approximations are necessary to simulate GC formation as a part of galaxy formation in the cosmic density field. Along this line, two semi-analytic approaches have been developed, each incorporating different physical inputs derived from numerical simulations. The first approach involves post-processing cosmological hydrodynamic simulations, or their zoom-in runs, with subgrid physics specific to GC formation. Depending on the resolution of the simulation used, the input conditions for GC formation can range from global properties of individual GMCs (e.g. Grudić et al., 2023) to the smoothed field of the gas properties (e.g. Kruijssen et al., 2011; Pfeffer et al., 2018; Kruijssen et al., 2019a). The second approach more closely resembles traditional semi-analytic models of galaxies based on dark matter halos and their merger trees. The global properties of individual galaxies are obtained from a semi-analytic model (El-Badry et al., 2019; De Lucia et al., 2023) or an empirical model of galaxy formation (Li & Gnedin, 2014; Choksi et al., 2018; Valenzuela et al., 2021), while the criteria for GC formation are based on some properties of the assembly histories of galaxies/halos, such as mergers of galaxies (El-Badry et al., 2019) and rates of halo mass assembly (Li & Gnedin, 2014; Choksi et al., 2018; Valenzuela et al., 2021; Chen & Gnedin, 2022). Both approaches have been successful in reproducing some observed properties of GCs. Despite the progress, significant knowledge gaps still exist in our understanding of the GC population. Indeed, our understanding of galaxy formation, in particular of star formation and feedback processes, is still incomplete, and so the large-scale conditions for the formation of GCs are still uncertain. In addition, small-scale processes driving the formation and disruption of GCs are also not well understood.
@@
Recent observations by the James Webb Space Telescope (JWST) have opened a new avenue to study galaxy formation in the early Universe. Recent observations indicated that early galaxies may differ significantly from their counterparts in the local Universe, suggesting that the processes driving galaxy formation at high $z$ may be different from those in the low-$z$ Universe. In particular, the efficiency of gas assumption by star formation seems to be enhanced significantly at high $z$ (Bouwens et al., 2023; Finkelstein et al., 2023; Xiao et al., 2023; Weibel et al., 2024; Wang et al., 2024), supermassive black holes (SMBHs) appear to be over-massive for their host galaxy mass (Maiolino et al., 2023; Pacucci et al., 2023; Matthee et al., 2024; Pacucci & Loeb, 2024), there seems to be an inversion of metallicity gradients in both the ISM and circumgalactic medium (CGM; Wang et al., 2022; Lin et al., 2023), and the total mass of bound star clusters in galaxies seems to be excessively large (Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024; Whitaker et al., 2025). Given that many GCs are old, these discoveries at high $z$ have important implications for understanding GC formation.
@@
Motivated partly by these discoveries, Mo et al. (2024, hereafter Paper-I) recently proposed a ‘two-phase’ model of galaxy formation that highlights the transition from an early phase of fast halo accretion to a later phase of slow accretion. A critical part of this model is the formation of self-gravitating and turbulent gas clouds (SGCs) during the fast phase. Such an SGC is expected to fragment subsequently and form a dynamically hot system of dense sub-clouds (SCs), each comparable to a GC in terms of both size and mass. This model is capable of reproducing numerous properties of galaxies observed over a large range of redshift. In particular, its predictions for the sub-cloud formation are directly related to the formation of GCs. As we will show in this paper, the model naturally provides conditions for two channels for the formation of GCs, and the GCs produced in the two channels have distinctive properties. Our model thus provides a unified cosmological framework to understand the formation of both galaxies and their GCs, allowing the model to be tested on spatial scales ranging from star clusters to large-scale structures, and on timescales from $z\approx 0$ to $z\approx 10$.
@@
This paper is organized as follows. In §2, we describe the context of galaxy formation within which GCs emerge. In §3, we elaborate on our model of GC formation. In §4, we describe numerical implementations of our model, present its predictions and compare them with observations and with other models. In §5, we summarize our results and discuss their implications. Throughout this paper, we adopt a flat $\Lambda$CDM cosmology with parameters derived from the Planck2015 results (Planck Collaboration et al., 2016): the density parameters $\Omega_{\rm M,0}=0.3089$, $\Omega_{\rm B,0}=0.0486$, and $\Omega_{\rm\Lambda,0}=0.6911$; the Hubble constant $H_{0}=100\,h\,{\rm km\,s^{-1}\,Mpc^{-1}}$, with $h=0.6774$; the Gaussian initial density field with a power spectrum $P(k)\propto k^{n}$ with $n=0.9667$; and an amplitude specified by $\sigma_{8}=0.8159$. To avoid confusion, we use ‘log’ to denote the base-10 logarithm, and ‘ln’ to denote the natural logarithm. We use the terms 1, 2, and 3-$\sigma$ ranges to denote those centralized at the median value and covering $68\%$, $95\%$, and $99.7\%$, respectively, of the probability mass for any probability distribution.
@@ 2.1 The assembly of dark matter halos
Our modeling of the halo population is based on a sample of halos and merger trees constructed from numerical simulations of the current $\Lambda$CDM cosmology, as described in §4. For each subhalo merger tree rooted in a ‘root subhalo’ (i.e. the subhalo without a descendant), we decompose it into a set of disjoint branches using the method described in Chen et al. (2023, see their §3.2.1). This is equivalent to traversing over the tree in a depth-first order starting from the root subhalo, and extracting each subsequent segment of the tree bracketed by a subhalo and its main leaf. For each branch, we find the last snapshot when it is a central subhalo, and define this snapshot as the ‘infall snapshot’. The assembly history at and before this snapshot is defined as the ‘central stage’ of the branch, while that after it is defined as the ‘satellite stage’.
@@
For each branch in the merger tree, we smooth its virial mass history, $M_{\rm v}(z)$, by a running Gaussian kernel with a width equal to the dynamical timescale, $t_{\rm dyn}=R_{\rm v}(z)/V_{\rm v}(z)=1/[10H(z)]$, where $R_{\rm v}(z)$ and $V_{\rm v}(z)$ are, respectively, the virial radius and virial velocity of the halo, and $H(z)$ is the Hubble parameter, all evaluated at the redshift $z$ in question. Smoothed versions of other virial quantities are defined using the smoothed $M_{\rm v}(z)$, while histories of other relevant halo quantities are smoothed using the same window as that used for $M_{\rm v}(z)$. The smoothing is required to ensure the numerical stability of our model implementation. Note that we do not aim to model the short-time-scale variations in the properties of halos, galaxies and star clusters, but intend to model the long-time-scale averages. The results of our model should thus be understood as an ‘effective’ description over a timescale of $\gtrsim t_{\rm dyn}$.
@@ 2.2 The collapse of the gas component
where the cosmic baryon number density $n_{\rm B,0}=3H_{0}^{2}f_{\rm B}\Omega_{\rm m}/(8\pi Gm_{\rm p})\approx 2.48\times 10^{-7}\,{\rm cm}^{-3}$, $\Delta_{\rm v}=200$, $f_{\rm gas}=0.04f_{\rm gas,0.04}$, $\mu=1.2$. The factor $f_{\rm str}$ characterizes the ‘streamness’ of the gas inflow. For spherically symmetric contraction, $f_{\rm str}=1$, and the typical value of $f_{\rm gas}$, $0.04$, is regulated by feedback (see Chen et al., 2024, hereafter Paper-II). For cold-mode streams, $1<f_{\rm str}\leq f_{\rm B}/f_{\rm gas}\approx 4$, as gas can penetrate into a galaxy along filaments and a larger fraction of gas can pile up around the halo center. Mandelker et al. (2018) and Dekel et al. (2023) estimated that the effective radius of streams is about $0.043R_{\rm v}$, similar to the contraction factor based on $f_{\rm gas}$. The SGC density given by Eq. (2.2) is much lower than the supernova-free threshold $n_{\rm snf}\approx 10^{3.5}\,{\rm cm}^{-3}$ required to form GCs (see §3.5) at $z=2$, and marginally touches the threshold at $z=9$. Thus, additional processes within SGCs are required for the formation of GCs in galaxies.